use the normal curve approximation to the binomial distribution.

two percent of hair dryers produced in a certain plant are defective. Estimate the probability that of 10,000 randomly selected hair dryers, at least 219 aredefective.
Answer #1

Mean is 200
Standard deviation is sqrt(.02*.98*10,000)= 14
z score is (218.5-200)/14= 1.32 (218.5 is for continuity correction)
Checking normal table probability of a z score of 1.32 or greater is .0934.

answered by: Marylyn